# Yuehao Bai

### v7

@@ 3 The Asymptotic Variance of Finely Stratified Experiments
In this section, we derive the asymptotic distribution of the method of moments estimator $\hat{\theta}_{n}$ when treatment is assigned by fine stratification over the baseline covariates $X^{(n)}$. Such assignment mechanisms use the covariates $X^{(n)}$ to group units with similar covariate values into blocks of fixed size, and then assign treatment completely at random within each block. In order to describe this assignment mechanism formally, we require some further notation to define the blocks of units. Let $\ell$ and $k$ be arbitrary positive integers with $\ell<k$ and set $\eta=\ell/k$. For simplicity, assume that $n$ is divisible by $k$. We then represent blocks of units using a partition of $\{1,\ldots,n\}$ given by
@@
with $|\lambda_{j}|=k$. Because of its possible dependence on $X^{(n)}$, $\{\lambda_{j}:1\leq j\leq n/k\}$ encompasses a variety of different ways of blocking the $n$ units according to the observed, baseline covariates. We note, however, that our framework is not intended to reflect settings in which the blocks themselves are randomly sampled from the population of interest (see, for instance, the discussion in Section 4.4 of pashley2021insights). Given such a partition, we assume that treatment status is assigned as described in the following assumption:
@@
$\{\pi(2j-1),\pi(2j)\}\text{ for }j=1,\ldots,n/2~,$
where $\pi=\pi_{n}(X^{(n)})$ is a permutation of $n$ elements. Because of its possible dependence on $X^{(n)}$, $\pi$ encompasses a broad variety of ways of pairing the $n$ units according to the observed, baseline covariates $X^{(n)}$. Given such a $\pi$, we assume that treatment status is assigned so that Assumption 3.1 holds and, conditional on $X^{(n)}$, $(A_{\pi(2j-1)},A_{\pi(2j)}),j=1,\ldots,n/2$ are i.i.d. and each uniformly distributed over the values in $\{(0,1),(1,0)\}$. For some examples of such an assignment mechanism being used in practice, see, for instance, angrist2009effects, banerjee2015miracle, and bruhn2016impact.
@@
Note that Assumption 3.1 generalizes matched pairs designs along two dimensions: first, it allows for treatment fractions other than $\eta=1/2$. Second, it allows for choices of $\ell$ and $k$ which are not relatively prime. For instance, if we set $(\ell,k)=(2,4)$, then $\eta=1/2$ as in matched pairs, but now the assignment mechanism blocks units into groups of size $4$ and assigns two units to treatment, two units to control. Although Theorem 3.1 below establishes that allowing for this level of flexibility has no effect on the asymptotic properties of our estimator, in our experience we have found that designs which employ these treatment “replicates” in each block can simplify the construction of variance estimators in practice; see Section 3.1 for details, and imbens2011experimental for an early discussion.
@@
bai2022inference and cytrynbaum2023designing discuss blocking algorithms that satisfy Assumption 3.2. When $X_{i}\in\mathbb{R}$ and $E_{Q}[X_{i}^{2}]<\infty$, a simple algorithm that satisfies Assumption 3.2 is simply to order units from smallest to largest and then block adjacent units into blocks of size $k$. In the case of matched pairs, if $\mathrm{dim}(X_{i})>1$ and $E_{Q}[\|X_{i}\|^{d}]<\infty$ for $d\geq\mathrm{dim}(X_{i})+1$, then Assumption 3.2 is satisfied by the nbpmatching algorithm in R that minimizes the sum of squared distances of $X$ within pairs. See Appendix A of bai2024inference-1 for details. Beyond the case of pairs, cytrynbaum2023designing demonstrates that the optimal blocking satisfies the following bound
@@
from which we can deduce that the rate of convergence depends on both the dimension of $X_{i}$ as well as the number of moments it possesses. Finally, we impose the following assumptions to derive the large-sample properties of $\hat{\theta}_{n}$. In what follows, when writing expectations and variances, we suppress the subscripts $P$ and $Q$ whenever doing so does not lead to confusion.
@@
Assumption 3.3(a) is a standard assumption to ensure the solution to (3) is “well separated.” It appears as a condition, for instance, in Theorem 5.9 in van_der_vaart1998asymptotic. Assumption 3.3(b) is a standard assumption used when deriving the properties of $Z$-estimators. See, for instance, Theorem 3.1 in newey1994large and Theorem 5.21 in van_der_vaart1998asymptotic. Because differentiability is imposed on their expectations instead of the moment functions themselves, the moment functions are allowed to be nonsmooth, as in Example 2.2. Assumption 3.3(c) requires the moment function to be mean-square continuous in $\theta$. Assumption 3.3(d) is a standard condition to guarantee the measurability of the supremum of a suitable class of functions. In particular, it allows us to define expectations of suprema without invoking outer expectations. See Example 2.3.4 in van_der_vaart1996weak for details. Assumption 3.3(e) is a standard assumption which guarantees the existence of an integrable envelope and allows us to invoke a uniform law of large numbers and a uniform central limit theorem (see, for instance, page 81 of van_der_vaart1996weak for a definition of a Donsker class). In particular, this assumption can be verified for Examples 2.12.5. Assumption 3.3(f) is a common assumption which simplifies some arguments when studying finely stratified designs, and ensures units that are close in terms of the baseline covariates are also close in terms of their moments. Note that Assumption 3.3(f) could be dropped following the approximation arguments in Lemma C.5 of cytrynbaum2023designing; see also Examples 4.4 and 4.9 for further discussion.
The following theorem establishes the asymptotic variance of the naïve method of moments estimator when the treatment assignment mechanism is finely stratified in the sense of satisfying Assumptions 3.13.2. Its proof relies on a crucial technical result in han2021complex, which allows us to compare the empirical process that depends on the treatment assignments with the empirical process that only depends on i.i.d. quantities. As a consequence of us leveraging this result, Assumption 3.3 is comparable to the typical assumptions imposed to study the properties of method of moments estimators using i.i.d. data.
@@
In order to make Theorem 3.1 useful for inference about $\theta_{0}$, we describe in Section 3.1 an estimator $\hat{\mathbb{V}}_{n}$ of $\mathbb{V}_{*}$. We now sketch an argument of the proof of Theorem 3.1 to highlight the fundamental role played by a “fast-balancing” property of finely stratified designs (see (14) below). In the proof of Theorem 3.1, we first establish that
@@
To further establish (9), it thus suffices to show that, under a finely stratified design,
@@
We call (14) the fast-balancing condition for the function $m(\cdot)$. Intuitively, the fast-balancing condition imposes that the experimental design should balance the treatment across covariate values at a rate which is faster than sampling variation. To see why (14) holds for a finely stratified design, let $\Omega(X_{i})=E[m(X_{i},1,R_{i}(1),\theta_{0})-m(X_{i},0,R_{i}(0),\theta_{0})|X_{i}]$ and first note that, by Assumption 2.1,
@@
which matches the asymptotic variance derived in bai2022inference for matched pairs. Theorem 3.1 however accommodates a much larger class of parameters, including those introduced in Examples 2.22.5.
@@
which is positive semidefinite. From this, we conclude that the asymptotic variance of the naive method of moments estimator $\hat{\theta}_{n}$ is lower in a finely stratified design compared to i.i.d. assignment. In Section 4.1, we will further show that $\mathbb{V}_{\ast}$ is the lowest possible asymptotic variance among regular estimators for $\theta_{0}$ in a large class of treatment assignment mechanisms, including both i.i.d. assignment and finely stratified designs. When $d_{\theta}=1$, we may express $\mathbb{V}-\mathbb{V}_{\ast}$ in terms of the “nonparametric $R^{2}$.” In particular, $\mathbb{V}-\mathbb{V}_{\ast}$ is proportional to $E[R_{g,X}^{2}(g_{i},X_{i})\operatorname{Var}[g_{i}]]$, where
@@
and $g_{i}=m(X_{i},1,R_{i}(1),\theta_{0})-m(X_{i},0,R_{i}(0),\theta_{0})$. The quantity in (19) measures how much of the variation in $g_{i}$ can be explained nonparametrically by $X_{i}$. See, for instance, chernozhukov2024long.
@@
Note that finely stratified designs include as a special case completely randomized designs, i.e., experiments in which a fixed proportion of the entire sample is assigned to treatment uniformly at random. To see this, consider, for instance, simply matching on an exogenously generated covariate. In this special case, we find from (18) that $\mathbb{V}_{\ast}=\mathbb{V}$ as defined in Section 2. In this way, we see that completely randomized experiments are asymptotically no more efficient than i.i.d. assignment.
@@ 3.1 Variance Estimation
In examples including Example 2.2 where $m$ is nonsmooth in $\theta$, $M$ may consist of components that require nonparametric estimators. See, for instance, jiang2021bootstrap.
@@
The estimator $\hat{\varsigma}_{n}(1,1)$ is constructed in one of two ways depending on the number of treated units in each block. If more than one unit in each block is treated, then we take the averages of all pairwise products of the treated units in each block, and average them across all blocks. We call this a “within block” estimator. If instead only one unit in each block is treated, then we take the product of two treated units in adjacent blocks. We call this a “between block” estimator, and note that similar constructions have been used previously in abadie2008estimation, bai2022inference, bai2024inference-1, and cytrynbaum2023designing. The estimator $\hat{\varsigma}_{n}(0,0)$ is constructed similarly. A natural estimator for $\mathbb{V}_{\ast}$ is then given by
@@
Note that for specific choices of $m(\cdot)$, $\hat{\mathbb{V}}_{n}$ recovers estimators which have been studied in prior work on inference in finely stratified experiments. For instance, in the case of matched pairs with $(\ell,k)=(1,2)$ and $m(\cdot)$ as in Example 2.1, so that $\theta_{0}$ is the ATE, $\hat{\mathbb{V}}_{n}$ exactly coincides with the estimator defined in equation (28) of bai2022inference.
@@
Note that given blocks which satisfy Assumption 3.2, it is always possible to re-order the blocks such that the pairs $\{(\lambda_{2j-1},\lambda_{2j})\}_{1\leq j\leq\lfloor n/2\rfloor}$ satisfy Assumption 3.4, as long as we maintain the sufficient condition that $E[\|X_{i}\|^{d}]<\infty$ for $d\geq\mathrm{dim}(X_{i})+1$. This property could be achieved, for instance, by applying the nbpmatching algorithm to the block-means $\{\bar{X}_{j}\}_{1\leq j\leq n/k}$, where $\bar{X}_{j}:=\frac{1}{k}\sum_{i\in\lambda_{j}}X_{i}$: see Lemma A.6 in cytrynbaum2023covariate for details.
@@
Assumption 3.5(a) is a mild uniform integrability condition for the envelope function of the moment function in an arbitrarily small neighborhood of $\theta_{0}$. Assumption 3.5(b) is a mild condition that requires the conditional expectation of the moment functions to satisfy a uniform law of large numbers. See page 81 of van_der_vaart1996weak for a definition of a Glivenko-Cantelli class of functions. Assumption 3.5(c) strengthens Assumption 3.3(f) to hold uniformly in an arbitrarily small neighborhood of $\theta_{0}$.
@@ 4 An Efficiency Bound and the Necessity of “Fast-Balancing”
In Section 4.1 we establish that $\mathbb{V}_{\ast}$ is the efficiency bound for a large class of experimental designs. As a consequence, we can conclude that finely stratified designs are asymptotically efficient “by design.” Building on this result, in Section 4.2 we establish that a necessary condition for achieving the bound $\mathbb{V}_{\ast}$ when estimating $\theta_{0}$ using the naïve method of moments estimator is that the experimental design be fast-balancing, in the sense of (14).
@@ 4.1 Efficiency Bound
An inspection of the asymptotic variance in (11) reveals that $\mathbb{V}_{\ast}$ in fact coincides with the classical efficiency bound for estimating $\theta_{0}$ with i.i.d. assignment. For example, the variance derived in (17) coincides with the efficiency bound derived in hahn1998role for estimating the ATE with a known marginal treatment probability $\eta$. Therefore, another way to interpret our result in Theorem 3.1 is that the standard i.i.d. efficiency bound can be attained by a naïve method of moments estimator under a finely stratified design. On the other hand, because treatment status is not independent in a finely stratified design, a natural follow-up question is whether or not the efficiency bound for estimating $\theta_{0}$ changes relative to what can be obtained under i.i.d. assignment once we allow for more general assignment mechanisms. In this section, we show that $\mathbb{V}_{*}$ continues to be the efficiency bound for the class of parameters introduced in Section 2, while allowing for a more general class of treatment assignment mechanisms. The main restriction on treatment assignment is given by Assumption 2.1, which requires the marginal treatment probability to be known and equal to $\eta$. As mentioned earlier and explained in Remark 4.4 below, it is possible to relax this requirement so that $\eta$ can be replaced by a known function $\eta(X_{i})$. For a discussion of how our efficiency bound compares with other results in the literature, see Remark 4.3.
@@
The treatment assignment mechanism is such that for any integrable function $\gamma:\mathbb{R}^{d_{x}}\to\mathbb{R}$,
@@
Let $S:\mathbf{R}^{d_{x}}\to\mathcal{S}=\{1,\ldots,\allowbreak|\mathcal{S}|\}$ be a function that maps the covariates into a fixed, finite set of discrete strata. We call such a stratification “coarse”, to distinguish it from “fine” stratification as defined in Section 3. Define $S_{i}=S(X_{i})$ and assume that treatment status is assigned so that
@@
This high-level assumption accommodates a large class of stratified assignment mechanisms, including stratified biased coin designs (efron1971forcing; wei1978adaptive), minimization methods (pocock1975sequential; hu2012asymptotic) and stratified block randomization (see zelen1974randomization, for an early discussion). It follows from Lemma C.4 in bugni2019inference that for any integrable function $\gamma(\cdot)$,
@@
ma2024new propose a family of covariate adaptive randomization procedures which assign treatment sequentially based on an imbalance metric defined by “feature maps” of (potentially continuous) covariates. It follows by Theorem 3.5 of their paper that Assumption 4.1 is satisfied under appropriate conditions.
@@
Suppose $n$ is even and we assign treatment using a finely stratified design with $(\ell,k)=(1,2)$. As discussed at the beginning of Section 3, such a design is also known as a matched pairs design. Assume that the pairing algorithm $\pi_{n}(X^{(n)})$ results in pairs that are close in the sense of Assumption 3.2. It then follows from the same argument used to establish (14) in the discussion following Theorem 3.1 that for any Lipschitz integrable function $\gamma(\cdot)$,
@@
By approximating integrable functions by Lipschitz integrable functions as in Lemma A.1 in hanneke2021universal, it can be shown that the convergence holds for any integrable function $\gamma(\cdot)$. Therefore, Assumption 4.1 is satisfied.
@@
Re-randomization is an assignment mechanism in which researchers specify a balance criterion for the covariates, and then repeatedly generate assignments using a completely randomized design until an assignment is found which achieves an acceptable covariate distribution according to the balance criterion. The properties of re-randomization procedures have been studied in li2017general; li2020rerandomization, li2018asymptotic; li2020rerandomization-1, and cytrynbaum2024finely. It follows from Corollary 3.7 in cytrynbaum2024finely that Assumption 4.1 holds for re-randomization designs, under appropriate assumptions.
We now present an efficiency bound for the parameter $\theta_{0}$ introduced in Section 2. Formally, we characterize the bound via a convolution theorem that applies to all regular estimators of the parameter $\theta_{0}$. Following the definition on page 365 of van_der_vaart1998asymptotic, by a regular estimator, we mean an estimator whose asymptotic distribution is invariant to “local” perturbations of the data generating process:
@@
where $P_{t/\sqrt{n},g}$ represents a “local” perturbation of the distribution $P$ along a “path” with “score” $g$. We leave the precise definition of regularity and related assumptions to Appendix A.2. In the paragraph following the statement of the theorem we provide some more details on the nature of our result.
@@
Suppose Assumptions 2.1, 3.3(b), and 4.1 hold, as well as Condition A.1 described in Appendix A.2. Further suppose $\mathbb{V}_{\ast}<\infty$. Let $\tilde{\theta}_{n}$ be any regular estimator of the parameter $\theta_{0}$ in the sense of (S.16) in Appendix A.2. Then,

### v8

@@ 3 The Asymptotic Variance of Highly Stratified Experiments
In this section, we derive the asymptotic distribution of the method of moments estimator $\hat{\theta}_{n}$ when treatment is assigned by a highly stratified design over the baseline covariates $X^{(n)}$. This assignment mechanism uses the covariates $X^{(n)}$ to group units with similar covariate values into blocks of fixed size, and then assigns treatment completely at random within each block. In order to describe the assignment mechanism formally, we require some further notation to define the blocks of units. Let $\ell$ and $k$ be arbitrary positive integers with $\ell<k$ and set $\eta=\ell/k$. Here, $k$ is the total number of units in each block and $\ell$ is the number of treated units in each block. For simplicity, assume that $n$ is divisible by $k$. We then represent blocks of units using a partition of $\{1,\ldots,n\}$ given by
@@
with $|\lambda_{j}|=k$. Because of its possible dependence on $X^{(n)}$, $\{\lambda_{j}:1\leq j\leq n/k\}$ encompasses a variety of different ways of blocking the $n$ units according to the observed, baseline covariates. We note, however, that our framework is not intended to reflect settings in which the blocks themselves are randomly sampled from the population of interest (see, for instance, the discussion in Section 4.4 of Pashley and Miratrix, 2021). Given such a partition, we assume that treatment status is assigned as described in the following assumption:
@@
$\lambda_{j}=\{\pi(2j-1),\pi(2j)\}\text{ for }j=1,\ldots,n/2~,$
where $\pi=\pi_{n}(X^{(n)})$ is a permutation of $n$ elements. Because of its possible dependence on $X^{(n)}$, $\pi$ encompasses a broad variety of ways of pairing the $n$ units according to the observed, baseline covariates $X^{(n)}$. Given such a $\pi$, we assume that treatment status is assigned so that Assumption 3.1 holds and, conditional on $X^{(n)}$, $(A_{\pi(2j-1)},A_{\pi(2j)}),j=1,\ldots,n/2$ are i.i.d. and each uniformly distributed over the values in $\{(0,1),(1,0)\}$. For some examples of such an assignment mechanism being used in practice, see, for instance, Angrist and Lavy (2009), Banerjee et al. (2015), and Bruhn et al. (2016).
@@
Note that Assumption 3.1 generalizes matched pairs designs along two dimensions: first, it allows for treatment fractions other than $\eta=1/2$. Second, it allows for choices of $\ell$ and $k$ which are not relatively prime. For instance, if we set $(\ell,k)=(2,4)$, then $\eta=1/2$ as in matched pairs, but now the assignment mechanism blocks units into groups of size $4$ and assigns two units to treatment, two units to control. Although Theorem 3.1 below establishes that allowing for this level of flexibility has no effect on the asymptotic properties of our estimator, in our experience we have found that designs which employ these treatment “replicates” in each block can simplify the construction of variance estimators in practice; see Section 3.1 for details, and Imbens (2011) for an early discussion.
@@
Bai et al. (2022) and Cytrynbaum (2023b) discuss blocking algorithms that satisfy Assumption 3.2. When $X_{i}\in\mathbf{R}$ and $E_{Q}[X_{i}^{2}]<\infty$, a simple algorithm that satisfies Assumption 3.2 is simply to order units from smallest to largest and then block adjacent units into blocks of size $k$. In the case of matched pairs, if $\mathrm{dim}(X_{i})>1$ and $E_{Q}[\|X_{i}\|^{d}]<\infty$ for $d\geq\mathrm{dim}(X_{i})+1$, then Assumption 3.2 is satisfied by the nbpmatching algorithm in R that minimizes the sum of squared distances of $X$ within pairs: see Appendix A of Bai et al. (2024b) for details. Beyond the case of pairs, Cytrynbaum (2023b) demonstrates that the optimal blocking satisfies the following bound
@@
from which we can deduce that the rate of convergence depends on both the dimension of $X_{i}$ as well as the number of moments it possesses. Note further that Assumption 3.2 can be satisfied even if $X_{i}$ contains discrete components, as may arise when summarizing a categorical variable numerically using, e.g., one-hot encoding.
Finally, we impose the following assumptions to derive the large-sample properties of $\hat{\theta}_{n}$. In what follows, when writing expectations and variances, we suppress the subscripts $P$ and $Q$ whenever doing so does not lead to confusion.
@@
Assumption 3.3(a) is a standard assumption to ensure the solution to (3) is “well separated.” It appears as a condition, for instance, in Theorem 5.9 in van der Vaart (1998). Assumption 3.3(b) is a standard assumption used when deriving the properties of $Z$-estimators. See, for instance, Theorem 3.1 in Newey and McFadden (1994) and Theorem 5.21 in van der Vaart (1998). Because differentiability is imposed on their expectations instead of the moment functions themselves, the moment functions are allowed to be nonsmooth, as in Example 2.2. Assumption 3.3(c) requires the moment function to be mean-square continuous in $\theta$. Assumption 3.3(d) is a standard condition to guarantee the measurability of the supremum of a suitable class of functions. In particular, it allows us to define expectations of suprema without invoking outer expectations. See Example 2.3.4 in van der Vaart and Wellner (1996) for details. Assumption 3.3(e) is a standard assumption which guarantees the existence of an integrable envelope and allows us to invoke a uniform law of large numbers and a uniform central limit theorem (see, for instance, page 81 of van der Vaart and Wellner (1996) for a definition of a Donsker class). In particular, this assumption can be verified for Examples 2.12.5. Assumption 3.3(f) is a common assumption which simplifies some arguments when studying highly stratified designs, and ensures units that are close in terms of the baseline covariates are also close in terms of their moments. Note that Assumption 3.3(f) could be dropped following the approximation arguments in Lemma C.5 of Cytrynbaum (2023b); see also Examples 4.4 and 4.9 for further discussion.
The following theorem establishes the asymptotic variance of the naïve method of moments estimator when the treatment assignment mechanism is highly stratified in the sense of satisfying Assumptions 3.13.2. Its proof relies on a crucial technical result in Han and Wellner (2021), which allows us to compare the empirical process that depends on the treatment assignments with the empirical process that only depends on i.i.d. quantities. As a consequence of us leveraging this result, Assumption 3.3 is comparable to the typical assumptions imposed to study the properties of method of moments estimators using i.i.d. data.
@@
In order to make Theorem 3.1 useful for inference about $\theta_{0}$, we describe in Section 3.1 an estimator $\hat{\mathbb{V}}_{n}$ of $\mathbb{V}_{*}$. We now sketch an argument of the proof of Theorem 3.1 to highlight the fundamental role played by a “fast-balancing” property of highly stratified designs (see (14) below). In the proof of Theorem 3.1, we first establish that
@@
To further establish (9), it thus suffices to show that, under a highly stratified design,
@@
We call (14) the fast-balancing condition for the function $m(\cdot)$. Intuitively, the fast-balancing condition imposes that the experimental design should balance the treatment across covariate values at a rate which is faster than sampling variation. To see why (14) holds for a highly stratified design, let $\Omega(X_{i})=E[m(X_{i},1,R_{i}(1),\theta_{0})-m(X_{i},0,R_{i}(0),\theta_{0})|X_{i}]$ and first note that, by Assumption 2.1,
@@
which matches the asymptotic variance derived in Bai et al. (2022) for matched pairs. Theorem 3.1 however accommodates a much larger class of parameters, including those introduced in Examples 2.22.5.
@@
which is positive semidefinite. From this, we conclude that the asymptotic variance of the naive method of moments estimator $\hat{\theta}_{n}$ is lower in a highly stratified design compared to i.i.d. assignment. In Section 4.1, we will further show that $\mathbb{V}_{\ast}$ is the lowest possible asymptotic variance among regular estimators for $\theta_{0}$ in a large class of treatment assignment mechanisms, including both i.i.d. assignment and highly stratified designs. When $d_{\theta}=1$, we may express $\mathbb{V}-\mathbb{V}_{\ast}$ in terms of the “nonparametric $R^{2}$.” In particular, $\mathbb{V}-\mathbb{V}_{\ast}$ is proportional to $E[R_{g,X}^{2}(g_{i},X_{i})\operatorname{Var}[g_{i}]]$, where
@@
and $g_{i}=m(X_{i},1,R_{i}(1),\theta_{0})-m(X_{i},0,R_{i}(0),\theta_{0})$. The quantity in (19) measures how much of the variation in $g_{i}$ can be explained nonparametrically by $X_{i}$. See, for instance, Chernozhukov et al. (2024).
@@
Note that highly stratified designs include as a special case completely randomized designs, i.e., experiments in which a fixed proportion of the entire sample is assigned to treatment uniformly at random. To see this, consider, for instance, simply matching on an exogenously generated covariate. In this special case, we find from (18) that $\mathbb{V}_{\ast}=\mathbb{V}$ as defined in Section 2. In this way, we see that completely randomized experiments are asymptotically no more efficient than i.i.d. assignment.
@@ 3.1 Variance Estimation
In examples including Example 2.2 where $m$ is nonsmooth in $\theta$, $M$ may consist of components that require nonparametric estimators. See, for instance, Jiang et al. (2021).
@@
The estimator $\hat{\varsigma}_{n}(1,1)$ is constructed in one of two ways depending on the number of treated units in each block. If more than one unit in each block is treated, then we take the averages of all pairwise products of the treated units in each block, and average them across all blocks. We call this a “within block” estimator. If instead only one unit in each block is treated, then we take the product of two treated units in adjacent blocks. We call this a “between block” estimator, and note that similar constructions have been used previously in Abadie and Imbens (2008), Bai et al. (2022), Bai et al. (2024b), and Cytrynbaum (2023b). The estimator $\hat{\varsigma}_{n}(0,0)$ is constructed similarly. A natural estimator for $\mathbb{V}_{\ast}$ is then given by
@@
Note that for specific choices of $m(\cdot)$, $\hat{\mathbb{V}}_{n}$ recovers estimators which have been studied in prior work on inference in highly stratified experiments. For instance, in the case of matched pairs with $(\ell,k)=(1,2)$ and $m(\cdot)$ as in Example 2.1, so that $\theta_{0}$ is the ATE, $\hat{\mathbb{V}}_{n}$ exactly coincides with the estimator defined in equation (28) of Bai et al. (2022).
@@
Note that given blocks which satisfy Assumption 3.2, it is always possible to re-order the blocks such that the pairs $\{(\lambda_{2j-1},\lambda_{2j})\}_{1\leq j\leq\lfloor n/2\rfloor}$ satisfy Assumption 3.4, as long as we maintain the sufficient condition that $E[\|X_{i}\|^{d}]<\infty$ for $d\geq\mathrm{dim}(X_{i})+1$. This property could be achieved, for instance, by applying the nbpmatching algorithm to the block-means $\{\bar{X}_{j}\}_{1\leq j\leq n/k}$, where $\bar{X}_{j}:=\frac{1}{k}\sum_{i\in\lambda_{j}}X_{i}$: see Lemma A.6 in Cytrynbaum (2023a) for details.
@@
Assumption 3.5(a) is a mild uniform integrability condition for the envelope function of the moment function in an arbitrarily small neighborhood of $\theta_{0}$. Assumption 3.5(b) is a mild condition that requires the conditional expectation of the moment functions to satisfy a uniform law of large numbers. See page 81 of van der Vaart and Wellner (1996) for a definition of a Glivenko-Cantelli class of functions. Assumption 3.5(c) strengthens Assumption 3.3(f) to hold uniformly in an arbitrarily small neighborhood of $\theta_{0}$.
@@ 4 An Efficiency Bound and the Necessity of “Fast-Balancing”
In Section 4.1 we establish that $\mathbb{V}_{\ast}$ is the efficiency bound for a large class of experimental designs. As a consequence, we can conclude that highly stratified designs are asymptotically efficient “by design.” Building on this result, in Section 4.2 we establish that a necessary condition for achieving the bound $\mathbb{V}_{\ast}$ when estimating $\theta_{0}$ using the naïve method of moments estimator is that the experimental design be fast-balancing, in the sense of (14).
@@ 4.1 Efficiency Bound
An inspection of the asymptotic variance in (11) reveals that $\mathbb{V}_{\ast}$ in fact coincides with the classical efficiency bound for estimating $\theta_{0}$ with i.i.d. assignment. For example, the variance derived in (17) coincides with the efficiency bound derived in Hahn (1998) for estimating the ATE with a known marginal treatment probability $\eta$. Therefore, another way to interpret our result in Theorem 3.1 is that the standard i.i.d. efficiency bound can be attained by a naïve method of moments estimator under a highly stratified design. On the other hand, because treatment status is not independent in a highly stratified design, a natural follow-up question is whether or not the efficiency bound for estimating $\theta_{0}$ changes relative to what can be obtained under i.i.d. assignment once we allow for more general assignment mechanisms. In this section, we show that $\mathbb{V}_{*}$ continues to be the efficiency bound for the class of parameters introduced in Section 2, while allowing for a more general class of treatment assignment mechanisms. The main restriction on treatment assignment is given by Assumption 2.1, which requires the marginal treatment probability to be known and equal to $\eta$. As mentioned earlier and explained in Remark 4.4 below, it is possible to relax this requirement so that $\eta$ can be replaced by a known function $\eta(X_{i})$. For a discussion of how our efficiency bound compares with other results in the literature, see Remark 4.3.
@@
The treatment assignment mechanism is such that for any integrable function $\gamma:\mathbf{R}^{d_{x}}\to\mathbf{R}$,
@@
Let $S:\mathbf{R}^{d_{x}}\to\mathcal{S}=\{1,\ldots,\allowbreak|\mathcal{S}|\}$ be a function that maps the covariates into a fixed, finite set of discrete strata. We call such a stratification “coarse”, to distinguish it from highly stratified designs as defined in Section 3. Define $S_{i}=S(X_{i})$ and assume that treatment status is assigned so that
@@
This high-level assumption accommodates a large class of stratified assignment mechanisms, including stratified biased coin designs (Efron, 1971; Wei, 1978), minimization methods (Pocock and Simon, 1975; Hu and Hu, 2012) and stratified block randomization (see Zelen, 1974, for an early discussion). It follows from Lemma C.4 in Bugni et al. (2019) that for any integrable function $\gamma(\cdot)$,
@@
Ma et al. (2024) propose a family of covariate adaptive randomization procedures which assign treatment sequentially based on an imbalance metric defined by “feature maps” of (potentially continuous) covariates. It follows by Theorem 3.5 of their paper that Assumption 4.1 is satisfied under appropriate conditions.
@@
Suppose $n$ is even and we assign treatment using a highly stratified design with $(\ell,k)=(1,2)$. As discussed at the beginning of Section 3, such a design is also known as a matched pairs design. Assume that the pairing algorithm $\pi_{n}(X^{(n)})$ results in pairs that are close in the sense of Assumption 3.2. It then follows from the same argument used to establish (14) in the discussion following Theorem 3.1 that for any Lipschitz integrable function $\gamma(\cdot)$,
@@
By approximating integrable functions by Lipschitz integrable functions as in Lemma A.1 in Hanneke et al. (2021), it can be shown that the convergence holds for any integrable function $\gamma(\cdot)$. Therefore, Assumption 4.1 is satisfied.
@@
Re-randomization is an assignment mechanism in which researchers specify a balance criterion for the covariates, and then repeatedly generate assignments using a completely randomized design until an assignment is found which achieves an acceptable covariate distribution according to the balance criterion. The properties of re-randomization procedures have been studied in Li and Ding (2017, 2020), Li et al. (2018, 2020), and Cytrynbaum (2024). It follows from Corollary 3.7 in Cytrynbaum (2024) that Assumption 4.1 holds for re-randomization designs, under appropriate assumptions.
We now present an efficiency bound for the parameter $\theta_{0}$ introduced in Section 2. Formally, we characterize the bound via a convolution theorem that applies to all regular estimators of the parameter $\theta_{0}$. Following the definition on page 365 of van der Vaart (1998), by a regular estimator, we mean an estimator whose asymptotic distribution is invariant to “local” perturbations of the data generating process:
@@
where $P_{t/\sqrt{n},g}$ represents a “local” perturbation of the distribution $P$ along a “path” with “score” $g$. We leave the precise definition of regularity and related assumptions to Supplement A.2. In the paragraph following the statement of the theorem we provide some more details on the nature of our result.
@@
Suppose Assumptions 2.1, 3.3(b), and 4.1 hold, as well as Condition A.1 described in Supplement A.2. Further suppose $\mathbb{V}_{\ast}<\infty$. Let $\tilde{\theta}_{n}$ be any regular estimator of the parameter $\theta_{0}$ in the sense of (S.16) in Supplement A.2. Then,